When looking through a drawer i found: 6 pairs of scissors, 4 buttons, 7 paper clips, 13 pieces of paper, 2 rubber bands, 1 marble, 8 un sharpened pencils, 3 ball point pens, 1 sharpie.
Find the following probabilities and label independent/dependent
P(a piece of paper)=
P(2 rubber bands) with replacement =
P(4 un sharpened pencils) without replacement =
P ($100 bill) =

Answer by stanbon(75887)

When looking through a drawer i found:
6 pairs of scissors,
4 buttons,
7 paper clips,
13 pieces of paper,
2 rubber bands,
1 marble,
8 unsharpened pencils,
3 ball point pens,
1 sharpie.
-----------------------------------
Find the following probabilities and label independent/dependent
P(a piece of paper)= = 13/45
---
P(2 rubber bands) with replacement = (2/45)^2
---------------------
P(4 unsharpened pencils) without replacement = 8C4/45C4 = 70/148995
------------------------
P ($100 bill) = 0
----------------------
Cheers,
Stan H.
----------------